Circular No. 8292

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004aa Further to IAUC 8276, A. Udalski reports the OGLE-III discovery

of an apparent supernova located at α = 4^{h}34^{h}02^{s}.90, δ = −67^{o}54′11′′.7 (equinox 2000.0), which is 1′′.0 west of the center of the apparent host galaxy. Approximate I-band OGLE-III-survey magnitudes for SN 2004aa: Feb. 14.0 UT, [22; 16.0, 20.2; 18.0, 19.6; 20.0, 19.1; 22.0, 18.7.

COMET C/2002 T7 (LINEAR) W. J. Altenhoff, F. Bertoldi, and K. Menten, Max-Planck-Institut für

Radioastronomie, Bonn; and C. Thum, Institut de Radio-Astronomie Millimétrique (IRAM), Granada, report the detection of 250-GHz continuum emission from comet C/2002 T7 with the MAMBO bolometer array at the IRAM 30-m telescope, yielding the following flux densities: Feb. 13.636 UT, S = 6.5 ± 1.1 mJy/beam; 16.697, 8.0 ± 1.3. These photometric observations with a beam of 11′′ register emission at the level of 1 mJy also in neighboring channels, suggesting that full-width-at-half-maximum emission extends to a distance of ∼ 20′′. Around the time of the comet’s closest approach to the earth in May, the flux density per beam is expected to increase by a factor 10, and the total flux density by > 100.

(1313) BERNA R. Behrend, Geneva Observatory, writes (on behalf of R. Roy, S.

Sposetti, N. Waelchli, D. Pray, N. Berger, C. Demeautis, D. Matter, R. Durkee, A. Klotz, D. Starkey, and V. Cotrez) that photometric observations obtained of the minor planet (1313) on eight nights during Feb. 6–16 show a lightcurve of amplitude 0.25 mag and suggest that this is a binary system with an orbital period of 1.061 ± 0.005 days, showing mutual eclipses and/or occultations near both rotational lightcurve minima with a duration of ∼ 0.09 day and depth ∼ 0.7 mag, the first being centered on Feb. 7.85 UT. The regular-appearing lightcurve is synchronized with the eclipse events, indicating that at least one of the two bodies is elongated and rotates synchronously with the orbital motion; the sharp eclipse/occultation events indicate that both components have approximately the same size. The maximum orbital separation observed from earth would be ∼ 0′′.03.

2004 DW

Corrigendum. On IAUC 8291, line 5, for albedo read phase-function slope

2004 February 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green